OpenStudy (anonymous):

What is the sin(-495)?

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

How'd you get that?

OpenStudy (anonymous):

495degrees equals 360o +135o 135o= 180o-45o

OpenStudy (anonymous):

sin(180)-sin(45)=0-1/2=-1/2